OpenStudy (anonymous):

Simplify (2t − 9) + (13t2 − 5) + (t + 1) 16t4 − 13 16t2 − 13 13t2 + 2t2 − 13 13t2 + 3t − 13

Parth (parthkohli):

Remove the brackets. They're of no use here. Then combine the like terms.

OpenStudy (anonymous):

(2t − 9) + (13t^2 − 5) + (t + 1) = 2t - 9 + 13t^2 - 5 + t + 1 = 13t^2 + 2t + t - 9 -5 + 1 = Can you simplify it?

OpenStudy (anonymous):

is it the third one

OpenStudy (anonymous):

Yes.